# CBMC-V3: A CNS-inspired Control Framework Towards Manipulation Agility with SNN

### v2

@@ III-C Thalamus module
As described in the previous sub-section, although we trained two neural network units under different end-effector load conditions to form the cerebellum module—thereby generating gravity compensation torques corresponding to two specific load cases—the cerebellum module alone remains insufficient for handling arbitrary load. To address this limitation, we developed the thalamus module. By utilizing torque sensor data at the robotic arm’s end-effector, the thalamus module network is able to “sense” the weight of the object, akin to human perception, and accordingly assign adaptive weights to the two gravity compensation torques produced by the cerebellum module’s motor patterns via reinforcement learning. This mechanism enables effective compensation for loads of varying mass and shape.
Figure 3: Structure of thalamus module.
The thalamus module contains five layers of neurons, as illustrated in Fig. 3. Its input $\bm{F}_{\text{ef}}\in\mathbb{R}^{6}$ and $\bm{T}_{\text{cb}}=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})\in\mathbb{R}^{7\times 2}$, corresponds to the six-dimensional force/torque sensor data measured at the end-effector, and the cerebellum module output, respectively. Its output, $\bm{w}=(w_{1},w_{2})^{T}\in\mathbb{R}^{2}$, represents the weight coefficients assigned to the two gravity compensation torques provided by the cerebellum module.
The input layer is fully connected to a hidden layer consisting of five LIF neurons, which in turn connects to an output layer of two non-spiking LIF neurons. The membrane potentials of the output neurons are then passed through a Softmax function to produce the two weight coefficients. The loss function designed for this network comprises two terms, minimizing both the joint position error and the joint velocity error, i.e.,
@@
$\tau_{m}$ Hidden layer: 5.0 $u_{\text{fire}}$ Hidden layer: 0.01
@@
After obtaining $\bm{w}$ from the output layer, it is then combined with data from the connecting fiber (linking the thalamus module and the cerebellum module) to generate a weighted output via matrix multiplication, yielding
$\begin{split}\delta\bm{u}_{\text{ft}}(t)&=(\delta u_{\text{ft,1}}(t),\delta u_{\text{ft,2}}(t),\cdots,\delta u_{\text{ft,7}}(t))^{T}\\ &=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})\cdot\bm{w}=w_{1}\bm{\tau}_{\text{cb,L}}+w_{2}\bm{\tau}_{\text{cb,H}}\end{split}$ (17)
This weighted result is not directly used as the output; instead, it is injected as an increment to the membrane potentials of seven non-spiking LIF neurons that serve as a filter layer. Specifically, the membrane potential of each filtering-layer neuron at time $t$ is given by
$u_{\text{ft},i}(t)$ $=u_{\text{ft},i}(t-1)-\frac{u_{\text{ft},i}(t-1)-u_{\text{reset}}}{\tau_{m}}$ (18)
$+\frac{\delta u_{\text{ft},i}(t)}{\tau_{m}},i=1,\cdots,7$
Finally we take the membrane potential of each filter-layer neuron at time $t$ as the gravity compensation torque, i.e., $\bm{\tau}_{g}=\bm{u}_{\text{ft}}(t)=(u_{\text{ft},1}(t),u_{\text{ft},2}(t),\cdots,u_{\text{ft},7}(t))$, which is then transmitted to the spinal cord module for further processing. The parameters of the neurons in each layer of this network, as well as the training parameters, are summarized in Table II. During trajectory tracking tasks, the learning procedure of the thalamus module follows Algorithm 4, which can be directly implemented as Sub-thread 3 in Algorithm 1.
@@
1:  initialize gravity compensation torque $\bm{\tau}_{g}$, weight coefficients $\bm{w}$, network loss lossloss, optimizer $optim(lr,mo)$
@@ III-D Brainstem module
As illustrated in Fig. 1c, the primary role of the brainstem module is to dynamically select appropriate feedback control coefficients during operation through learning, thereby compensating for the nonlinear dynamics of the robotic arm and the coupling effects among joints. The network is implemented as a three-layer fully connected SNN. Its input variable $\bm{q}_{d},\bm{\dot{q}}_{d},\bm{q},\bm{\dot{q}}]\in\mathbb{R}^{28}$ is a concatenated vector consisting of the desired joint positions and velocities, as well as the actual joint positions and velocities. The output $\bm{K}=(k_{p,1},k_{p,2},\cdots,k_{p,7},k_{v,1},k_{v,2},\cdots,k_{v,7})\in\mathbb{R}^{14}$ corresponds to the proportional and derivative coefficients for each of the seven joints.
@@
In this formulation, the first term minimizes the position error and the second term minimizes the velocity error, ensuring accurate trajectory tracking, while the third term penalizes torque usage to reduce joint energy consumption and prevent abrupt fluctuations in joint torque. Through this design, the network is able to learn feedback control coefficients that balance trajectory-tracking accuracy with reduced joint wear and energy cost.
Similar to the cerebellum module, the brainstem module adopts rate–based encoding with a time window length of $T$, and the approximate gradient function is chosen as the sigmoid function.
@@
1:  initialize coefficients $\bm{K}_{V},\bm{K}_{P}$, network loss lossloss, optimizer $optim(lr,mo)$
@@ III-E Spinal cord module
The primary function of the spinal cord module is to integrate proprioceptive information with the outputs of the aforementioned higher-level motor control modules, thereby generating real-time motion control commands. Its structural design is illustrated in Fig. 1c. For high-performance robotic manipulators, torque control typically requires the control algorithm to achieve frequencies of 1000 Hz, 1200 Hz, or even higher. Consequently, the spinal cord module must exhibit very high computational efficiency.
In our design, the input variables $\bm{q}_{d},\bm{\dot{q}}_{d},\bm{q},\bm{\dot{q}}]\in\mathbb{R}^{28}$ consist of the planned positions and velocities as well as the actual positions and velocities of all joints, while the output $\bm{\tau}\in\mathbb{R}^{7}$ represents the control torque for each joint. After entering the spinal cord module, the planned trajectory and proprioceptive feedback are first encoded into discrete spike trains, with each joint being processed independently. Specifically, the planned position and velocity signals $q_{d,i},\dot{q}_{d,i},i=1,\cdots,7$ serve to excite the input fibers, whereas the actual proprioceptive feedback $q_{i},\dot{q}_{i},i=1,\cdots,7$ inhibits the input fibers. Mathematically, this process is equivalent to computing their difference.
@@
$s_{q,i,j}=\begin{cases}0,&-u_{j}<\frac{q_{d,i}-q_{i}}{\Delta q_{i}}<u_{j},j=1,\cdots,n\\ 1,&\frac{q_{d,i}-q_{i}}{\Delta q_{i}}\geq u_{j},j=1,\cdots,\frac{n}{2}\\ -1,&-\frac{q_{d,i}-q_{i}}{\Delta q_{i}}\geq u_{j},j=\frac{n}{2},\cdots,n\end{cases}$ (22)
$s_{\dot{q},i,j}=\begin{cases}0,&-u_{j}<\frac{\dot{q}_{d,i}-\dot{q}_{i}}{\Delta\dot{q}_{i}}<u_{j},j=1,\cdots,n\\ 1,&\frac{\dot{q}_{d,i}-\dot{q}_{i}}{\Delta\dot{q}_{i}}\geq u_{j},j=1,\cdots,\frac{n}{2}\\ -1,&-\frac{\dot{q}_{d,i}-\dot{q}_{i}}{\Delta\dot{q}_{i}}\geq u_{j},j=\frac{n}{2},\cdots,n\end{cases}$ (23)
Here, $u_{j}\sim U(0,1),j=1,\cdots,n$ denotes $n$ independent uniformly distributed random variables, ${s_{q,i,j}},i=1,\cdots,7,j=1,\cdots,n$ represents the spike trains encoding the position information of each joint at the current time step, and ${s_{\dot{q},i,j}},i=1,\cdots,7,j=1,\cdots,n$ denotes the spike trains encoding the velocity information. The terms $\Delta q_{i}$ and $\Delta\dot{q}_{i},i=1,\cdots,7$ correspond to the maximum allowable position and velocity errors for each joint, respectively.
From an intuitive perspective, each input fiber is connected to a group of $n$ submodules, where the first half $\frac{n}{2}$ are responsible for processing positive inputs and the second half $\frac{n}{2}$ handle negative inputs. This structure corresponds to the agonist–antagonist pairing characteristic of human muscles. Each submodule generates a uniformly distributed random number in the range $[0,1]$. When the output of the input fiber is positive, the first $\frac{n}{2}$ neurons are encoded sequentially: if the absolute value of the input exceeds the random number, the encoding result is set to $1$; otherwise, it is 0. Conversely, when the output of the input fiber is negative, the second $\frac{n}{2}$ neurons are encoded sequentially: if the absolute value of the input exceeds the random number, the encoding result is set to 1-1; otherwise, it is 0. This encoding scheme can be implemented in hardware using simple logic gate circuits, thereby achieving very high computational efficiency.
The encoding precision is influenced by three parameters: $\Delta q_{i}$, $\Delta\dot{q}_{i}$, $i=1,\cdots,7$, and $n$. Smaller values of $\Delta q_{i}$ and $\Delta\dot{q}_{i}$ lead to higher encoding precision, but reduce the range of representable errors; a larger $n$ increases encoding precision but reduces computational efficiency.
Following this encoding process, two sets of discrete spike trains are generated—one representing position information and the other representing velocity information. These spike trains are then transmitted to stimulate the activity of non-spiking LIF neurons in the output layer, with stimulation strength modulated by the feedback control gains $k_{p,i}$ and $k_{v,i}$, $i=1,\cdots,7$ provided by the brainstem module, as illustrated by the blue and red dashed lines, respectively. Accordingly, at each time step, the membrane potential increments of the output non-spiking LIF neurons can be expressed as
@@
Here, $\bm{\tau}_{g}=(\tau_{g,1},\tau_{g,2},\cdots,\tau_{g,7})$ denotes the gravity compensation torque generated by the thalamus module. Consequently, the membrane potentials at each time step can be written as
$u_{\text{sp},i}(t)$ $=u_{\text{sp},i}(t-1)-\frac{u_{\text{sp},i}(t-1)-u_{\text{reset}}}{\tau_{m}}$ (25)
$+\frac{\delta u_{\text{sp},i}(t)}{\tau_{m}},i=1,\cdots,7$
Finally, similar to the thalamus module, we take $\bm{\tau}=\bm{u}_{\text{sp}}=(u_{\text{sp},1}(t),u_{\text{sp},2}(t),\cdots,u_{\text{sp},7}(t))$ as the final control torque, which will be sent in real time to the actuators. The neuronal parameters of the spinal cord module are summarized in Table IV.
@@
$\tau_{m}$ 2.0 n 100
Through this mechanism, the output torque at each time step incorporates both the current information and the historical information over a preceding time window, thereby making it more suitable than ANN for time-continuous tasks such as robotic control. During trajectory tracking tasks, the operational process of the spinal cord module is illustrated in Algorithm 6, which can be implemented as Sub-thread 1 within Algorithm1.
@@
IV Simulation analysis
To evaluate the rationality and effectiveness of the proposed framework, we conduct three simulation experiments that demonstrate contributions (1)–(3). In section IV-A and IV-B, we validate the design of the modules and in section IV-C, the design of the framework.
The algorithm is implemented in Python (version 3.12.4), and all neural networks are constructed using PyTorch [paszke2019pytorch] (version 2.4.0). The spiking neurons involved in the algorithm are implemented with the SNN learning framework SpikingJelly [doi:10.1126/sciadv.adi1480] (version 0.0.0.0.14). The simulation environment employed is PyBullet [coumans2021] (version 3.2.6), an open-source robotics simulation tool whose underlying physics engine is based on the open-source Bullet Physics SDK. The controlled robotic arm consists of seven actuated joints, seven links and a six-dimensional torque sensor mounted at the flange of the end-effector. The control cycle of each layer is set as $T_{1}=1$ms, $T_{2}=10$ms, $T_{3}=50$ms. The simulation time step is set to 1 ms, and the control frequency is configured to 1 kHz.
In all simulation experiments presented in this section, trajectory tracking accuracy is evaluated using the root-mean-square error (RMSE) of the joint positions. The RMSE at each time step is defined as
@@ IV Simulation analysis
$\text{RMSE}=\frac{1}{N}\sum_{t=1}^{N}\text{RMSE}(t)$ (27)
@@
Three types of trajectories are employed in the simulations: horizontal circle, inclined circle, and spatial figure-eight. The first two trajectories have already been defined in section III-B, while the third trajectory is defined as
spatial figure-eight: $\begin{cases}x=x+0.5R\sin(\frac{4\pi t}{T})\\ y=y+R\cos(\frac{2\pi t}{T})\\ z=z+z\sin(\frac{2\pi t}{T})\end{cases}$ (28)
where $x=0.61,y=0,z=0.3,R=0.14,T=3,z=0.08$.
Figure 4: Average feedback control coefficients of each joint across 3 trajectory tracking tasks.
IV-A Spinal cord and brainstem module validation
Physiological studies have demonstrated that joint stiffness and damping in the human arm exhibit an inverse correlation [bennett1992time]. Specifically, stiffness decreases when a joint begins to move and increases when at rest, while damping decreases as the joint approaches a target and increases at the onset of motion. In the proposed framework, the feedback control coefficients $\bm{K}_{V}$ and $\bm{K}_{P}$ generated by the brainstem module correspond precisely to damping and stiffness in feedback control. Fig. 4 illustrates the variations of average $k_{p}$ (blue solid lines) and $k_{v}$ (red dashed lines) for each joint across three trajectories. It can be observed that the two parameters generally exhibit opposite trends, as exemplified by joints 3–6 during 3–5 s and joints 4 and 6 during 1–2 s. This finding indicates that the brainstem module, through reinforcement learning, adjusts joint stiffness and damping in the spinal module in a manner that exhibits biomimetic characteristics, consistent with our design objectives. We also note that this feature is less pronounced in joints closer to the robot’s base, such as joints 1 and 2, likely due to stronger coupling effects from adjacent joints.

### v3

@@ III-C Thalamus module
As described in the previous sub-section, although we trained two neural networks to form the cerebellum module—thereby generating gravity compensation torques corresponding to two specific load masses—the cerebellum module alone remains insufficient for handling arbitrary load. To address this limitation, we developed the thalamus module. By utilizing torque sensor data from the robotic arm’s end-effector, the thalamus module network is able to “sense” the mass of the object, akin to human perception, and accordingly assign adaptive weights to the two gravity compensation torques produced by the cerebellum module’s motor patterns via reinforcement learning.
The thalamus module contains five layers of neurons, as illustrated in Fig. 3. Its input $\bm{F}_{\text{ef}}\in\mathbb{R}^{6}$ and $\bm{T}_{\text{cb}}=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})\in\mathbb{R}^{7\times 2}$, correspond to the six-dimensional force/torque sensor data measured at the end-effector, and the cerebellum module output, respectively. Its output, $\bm{\tau}_{g}\in\mathbb{R}^{7}$, represents the gravity compensation torque for arbitrary load mass within a certain range.
Figure 3: Structure of the thalamus module.
The input layer is fully connected to a hidden layer consisting of five LIF neurons, which in turn connects to an output layer of two non-spiking LIF neurons. The membrane potential of the output neurons are then passed through a Softmax function to produce the two weight coefficients. The network is trained via reinforcement learning, and the loss function designed for this network comprises two terms, minimizing both the joint position error and the joint velocity error, i.e.,
@@
$\tau_{\text{mb}}$ Hidden layer: 5.0 $u_{\text{fire}}$ Hidden layer: 0.01
@@
After obtaining $\bm{w}$ from the output layer, it is then combined with data from the connecting fiber (linking the thalamus module and the cerebellum module) to generate a weighted output, yielding
$\begin{split}\delta\bm{u}_{\text{ft}}(t)&=(\delta u_{\text{ft,1}}(t),\delta u_{\text{ft,2}}(t),\cdots,\delta u_{\text{ft,7}}(t))^{\top}\\ &=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})\cdot\bm{w}=w_{1}\bm{\tau}_{\text{cb,L}}+w_{2}\bm{\tau}_{\text{cb,H}}\end{split}$ (17)
This would then be injected as an increment to the membrane potentials of seven non-spiking LIF neurons that serve as a filter layer. Specifically, the membrane potential of each filtering-layer neuron at time $t$ is given by
$u_{\text{ft},i}(t)$ $=u_{\text{ft},i}(t-1)-\frac{u_{\text{ft},i}(t-1)-u_{\text{reset}}}{\tau_{\text{mb}}}$ (18)
$+\frac{\delta u_{\text{ft},i}(t)}{\tau_{\text{mb}}},i=1,\cdots,7$
Finally we take the membrane potential of each filter-layer neuron at time $t$ as the gravity compensation torque, i.e., $\bm{\tau}_{g}=\bm{u}_{\text{ft}}(t)=(u_{\text{ft},1}(t),u_{\text{ft},2}(t),\cdots,u_{\text{ft},7}(t))^{\top}$. The parameters of the neurons in each layer of this network, as well as the training parameters, are summarized in Table II. During trajectory tracking tasks, the learning procedure of the thalamus module follows Algorithm 4, which can be directly implemented as Sub-thread 3 in Algorithm 1.
@@
1:  initialize gravity compensation torque $\bm{\tau}_{g}$, weight coefficients $\bm{w}$, learning rate lrlr, momentum momo, network loss lossloss, optimizer $optim(lr,mo)$
@@ III-D Brainstem module
As stated in section II-A4, the brainstem module is to provide appropriate feedback control coefficients during operation to compensate for the nonlinear dynamics of the robotic arm and the coupling effects among joints. Its network is implemented as a three-layer fully connected SNN, as illustrated in Fig. 1c. The input variable $\bm{q}_{d},\bm{\dot{q}}_{d},\bm{q},\bm{\dot{q}}]\in\mathbb{R}^{28}$ is a concatenated vector consisting of the desired joint positions and velocities, as well as the actual joint positions and velocities. The output $\bm{K}_{P}=diag(k_{p,1},k_{p,2},\cdots,k_{p,7})$, $\bm{K}_{V}=diag(k_{v,1},k_{v,2},\cdots,k_{v,7})$ corresponds to the proportional and derivative coefficients for each of the seven joints.
@@
In this formulation, the first term and second term minimize position and tracking error to ensure accurate trajectory tracking, while the third term penalizes torque usage to reduce joint energy consumption.
Similar to the cerebellum module, the brainstem module adopts rate–based encoding with a time window length of $T$, and the surrogate gradient function is chosen as the sigmoid function.
@@
1:  initialize coefficients $\bm{K}_{V},\bm{K}_{P}$, learning rate lrlr, momentum momo, network loss lossloss, optimizer $optim(lr,mo)$
@@ III-E Spinal cord module
According to section II-A5, the spinal cord module is designed to integrate proprioceptive information and the outputs of the aforementioned modules to generate real-time motor control commands. For high-performance robotic arms, torque control typically requires the control algorithm to achieve frequencies of 1 kHz or even higher. Consequently, the spinal cord module must exhibit very high computational efficiency.
The structure of this module is illustrated in Fig. 1c. The input variables $\bm{q}_{d},\bm{\dot{q}}_{d},\bm{q},\bm{\dot{q}}]\in\mathbb{R}^{28}$ consist of the planned positions and velocities as well as the actual positions and velocities of seven joints, while the output $\bm{\tau}\in\mathbb{R}^{7}$ represents the control torque for each joint. After entering the spinal cord module, the planned trajectory and proprioceptive feedback are first encoded into discrete spike trains, with each joint being processed independently. Specifically, the planned position and velocity signals $q_{d,i},\dot{q}_{d,i},i=1,\cdots,7$ excite the input fibers, whereas the actual proprioceptive feedback $q_{i},\dot{q}_{i},i=1,\cdots,7$ inhibit the input fibers. Mathematically, this process is equivalent to computing their difference.
@@
$s_{q,i,j}=\begin{cases}0,&-u_{j}<\frac{q_{d,i}-q_{i}}{\Delta q_{i}}<u_{j},j=1,\cdots,k\\ 1,&\frac{q_{d,i}-q_{i}}{\Delta q_{i}}\geq u_{j},j=1,\cdots,\frac{k}{2}\\ -1,&-\frac{q_{d,i}-q_{i}}{\Delta q_{i}}\geq u_{j},j=\frac{k}{2},\cdots,k\end{cases}$ (22)
$s_{\dot{q},i,j}=\begin{cases}0,&-u_{j}<\frac{\dot{q}_{d,i}-\dot{q}_{i}}{\Delta\dot{q}_{i}}<u_{j},j=1,\cdots,k\\ 1,&\frac{\dot{q}_{d,i}-\dot{q}_{i}}{\Delta\dot{q}_{i}}\geq u_{j},j=1,\cdots,\frac{k}{2}\\ -1,&-\frac{\dot{q}_{d,i}-\dot{q}_{i}}{\Delta\dot{q}_{i}}\geq u_{j},j=\frac{k}{2},\cdots,k\end{cases}$ (23)
Here, $u_{j}\sim U(0,1),j=1,\cdots,k$ denotes $k$ independent uniformly distributed random variables, ${s_{q,i,j}},{s_{\dot{q},i,j}},i=1,\cdots,7,j=1,\cdots,k$ denote the spike trains encoding the position and velocity information of each joint at each time step, respectively. The terms $\Delta q_{i}$ and $\Delta\dot{q}_{i},i=1,\cdots,7$ correspond to the maximum position and velocity errors for each joint, respectively.
From an intuitive perspective, each input fiber is connected to a group of $k$ neurons, where the first half $\frac{k}{2}$ are responsible for processing positive inputs and the second half $\frac{k}{2}$ handle negative inputs. This structure corresponds to the agonist–antagonist pairing characteristic of human muscles. Each neuron generates a uniformly distributed random number in the range of $[0,1]$. When the output of the input fiber is positive, the first $\frac{k}{2}$ neurons are activated: if the absolute value of each neuron’s input exceeds the random number, the encoding result is set to $1$; otherwise, it is 0. Conversely, when the output of the input fiber is negative, the second $\frac{k}{2}$ neurons are activated: if the absolute value of each neuron’s input exceeds the random number, the encoding result is set to 1-1; otherwise, it is 0. The encoding effect is influenced by three parameters: $\Delta q_{i}$, $\Delta\dot{q}_{i}$, $i=1,\cdots,7$, and $k$. Smaller values of $\Delta q_{i}$ and $\Delta\dot{q}_{i}$ lead to higher encoding precision, but reduce the dynamic range; a larger $k$ increases encoding precision but reduces computational efficiency.
Following this encoding process, two sets of spike trains are generated—one representing position information and the other representing velocity information. These spike trains are then transmitted to stimulate the activity of the non-spiking LIF neuron for each joint, with stimulation strength modulated by the feedback control coefficients $k_{p,i}$ and $k_{v,i}$, $i=1,\cdots,7$ provided by the brainstem module. Accordingly, at each time step, the membrane potential increment of the output non-spiking LIF neurons can be expressed as
@@
Here, $\bm{\tau}_{g}=(\tau_{g,1},\tau_{g,2},\cdots,\tau_{g,7})^{\top}$ denotes the gravity compensation torque generated by the thalamus module. Consequently, the membrane potentials at each time step can be written as
$u_{\text{sp},i}(t)$ $=u_{\text{sp},i}(t-1)-\frac{u_{\text{sp},i}(t-1)-u_{\text{reset}}}{\tau_{\text{mb}}}$ (25)
$+\frac{\delta u_{\text{sp},i}(t)}{\tau_{\text{mb}}},i=1,\cdots,7$
Finally, similar to the thalamus module, we take $\bm{\tau}=\bm{u}_{\text{sp}}=(u_{\text{sp},1}(t),u_{\text{sp},2}(t),\cdots,u_{\text{sp},7}(t))^{\top}$ as the control torque to be sent in real time to the actuators. The neuronal and encoding parameters of the spinal cord module are summarized in Table IV.
@@
$\tau_{\text{mb}}$ 2.0 k 100
During trajectory tracking tasks, the operational process of the spinal cord module is illustrated in Algorithm 6, which can be implemented as Sub-thread 1 within Algorithm1.
@@
IV Simulation results
In this section, we conduct simulation experiments to validate that the module designs proposed in Section III satisfy the requirements outlined in Section II, and to evaluate the resilience of the proposed framework.
The framework is implemented in Python (version 3.12.4), and all modules are constructed using PyTorch [34] (version 2.4.0). The spiking neurons involved in the framework are implemented with the SNN learning framework SpikingJelly [14] (version 0.0.0.0.14). The simulation environment employed is PyBullet [11] (version 3.2.6), an open-source robotics simulation tool whose underlying physics engine is based on the open-source Bullet Physics SDK. The controlled robotic arm is a Flexiv Rizon 4s consisting of seven actuated joints, seven links and a six-dimensional torque sensor mounted at the flange of the end-effector. The framework runs on a multi-core CPU, and the control cycle of the first-order, second-order and third-order level is set as $T_{1}=1$ms, $T_{2}=10$ms, $T_{3}=50$ms, respectively. The simulation time step is set to 1 ms, and the control frequency is configured to 1 kHz.
In all simulation experiments presented in this section, trajectory tracking accuracy is evaluated using the root mean square error (RMSE) of the joint positions. The RMSE at each time step is defined as
@@ IV Simulation results
$\text{RMSE}=\frac{1}{L}\sum_{t=1}^{L}\text{RMSE}(t)$ (27)
@@
where $L$ is the trajectory length. Three types of trajectories are employed in the simulations: horizontal circle, inclined circle, and spatial figure-eight. The first two trajectories have already been defined in section III-B, while the third trajectory is defined as
spatial figure-eight: $\begin{cases}x=x_{0}+0.5R_{e}\sin(\frac{4\pi t}{T_{e}})\\ y=y_{0}+R_{e}\cos(\frac{2\pi t}{T_{e}})\\ z=z_{0}+0.08\sin(\frac{2\pi t}{T_{e}})\end{cases}$ (28)
where $x_{0}=0.61,y_{0}=0,z_{0}=0.3,R_{e}=0.14,T_{e}=3$.
IV-A Module design validation
Fig. 4 illustrates the performance of the cerebellum and thalamus modules. As shown in Fig. 4a, when only the light motor pattern is used, the tracking error increases with increasing load mass, whereas when only the heavy motor pattern is used, the tracking error decreases as the load mass increases. This indicates that the cerebellum module effectively encodes two distinct motor patterns: the light pattern corresponds to a load of approximately 1 kg, while the heavy pattern corresponds to a load of approximately 3.5 kg. However, neither pattern alone can accommodate arbitrary load masses within the range of 1 kg to 3.5 kg.
Owing to its online learning capability, the thalamus module can adaptively adjust the weighting coefficients assigned to these two motor patterns. As illustrated in Fig. 4b, the thalamus module assigns a smaller weight to the light pattern and a larger weight to the heavy pattern as the load mass increases. Consequently, the weighted combination of the two patterns achieves consistently good performance across the entire load range from 1 kg to 3.5 kg. This effect is evident in Fig. 4a, where the black curve remains below the red and blue curves.
This adaptive mechanism mirrors human motor control, in which movements are generated through linear combinations of motor primitives [38], thereby highlighting the biomimetic nature of the cerebellum and thalamus modules. Overall, these results demonstrate that the designs of both modules satisfy the proposed requirements.